PID# Name Class day/time 8 9 1= 5$ FIND P 4 40 5 50 Flow -P 20 30

the rows above is EOY and Cash flow
Answer #1

ANSWER:

I = 5%

P = Cash flow in year 2(p/f,i,n) + Cash flow in year 3(p/f,i,n) + Cash flow in year 4(p/f,i,n) + Cash flow in year 5(p/f,i,n) + Cash flow in year 6(p/f,i,n) + Cash flow in year 7(p/f,i,n) + Cash flow in year 8(p/f,i,n) + Cash flow in year 9(p/f,i,n)

P = 20(P/F,5%,2) + 30(P/F,5%,3) + 40(P/F,5%,4) + 50(P/F,5%,5) - 5(P/F,5%,6) - 5(P/F,5%,7) - 5(P/F,5%,8) - 5(P/F,5%,9)

P = 20 * 0.907 + 30 * 0.8638 + 40 * 0.8227 + 50 * 0.7835 - 5 * 0.7462 - 5 * 0.7107 - 5 * 0.6768 - 5 * 0.6446

P = 18.14 + 25.914 + 32.908 + 39.175 - 3.731 - 3.55 - 3.38 - 3.22

P = 102.24

SO THE VALUE OF P IS $102.24